An online medical advice company just completed an IPO with an Investment bank on a firm-commitment basis. The firm issued five million shares of common stock, and the underwriting fees were $2.20 per share. The offering price was $27.30 per share. 


What were the total proceeds from the common-stock sale? 

Total proceeds _______ 


How much money did the company receive? 

Net proceeds to firm _______ 


How much money did the investment bank receive in fees? 

Underwriting spread _______ 
Answer #1
Total proceeds = Shares of common stock issued * Offering price per share = 5000000 * 27.30 136500000
Net proceeds to firm = Shares of common stock issued * ( Offering price per share - Underwriting fees per share ) = 5000000 * ( 27.30 - 2.20 ) 125500000
Underwriting spread = Shares of common stock issued * Underwriting fees per share = 5000000 * 2.20 11000000